Chemist must mix 30 % insecticide soln with 50% insecticide soln. (same insecticide) to obtain 200 L of a 42 % insecticide soln. Please help with 2 stage linear system of equation and demonstrate resolution. THX
Answer by checkley71(8403)

.5X+.3(200-X)=.42*200
.5X+60-.3X=84
.2X=84-60
.2X=24
X=24/.2
X=120 LITERS OF 50% IS USED.
200=120=80 LITERS OF 30% IS USED.
PROOF:
.5*120+.3*80=84
60+24=84
84=84